OpenStudy (anonymous):

-17 – x = -2(3x + 1)

OpenStudy (anonymous):

You need to distribute the -2 to each of the terms in the parentheses. Then you can solve it in the usual way.

OpenStudy (anonymous):

i did that polPak and heres what i came up with

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

Nope. What was the equation after you do the distributing?

OpenStudy (anonymous):

-17 - 6x +6X -2-X

OpenStudy (anonymous):

2(3X +1) -17- X= 6X + 2

OpenStudy (anonymous):

You forgot that it's a -2, not 2 you're distributing.

OpenStudy (anonymous):

\[-2(3x + 1) = -2(3x) -2(1) = -6x - 2\] So you should have \[17 - x = -6x -2\]

OpenStudy (anonymous):

oh so the 6 becomes -6x

OpenStudy (anonymous):

And the 2 becomes -2.

OpenStudy (anonymous):

Err My equation is slightly wrong.. You should have \[-17 -x = -6x -2\]

OpenStudy (anonymous):

-17 - x = -6 - 2

OpenStudy (anonymous):

Don't lose track of the x on the -6. It's important

OpenStudy (anonymous):

then now i can solve for x so the its -17x -6 - 2

OpenStudy (anonymous):

No. -17 - x is not -17x and you can't just put them together like that.

OpenStudy (anonymous):

Ok look at it like this.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You have this equation: \[-17 - x = -6x -2\] That means that \(-17 - x\) IS EXACTLY THE SAME AS \(-6x -2\) So if they are the same then isn't \((-6x -2) + 2\) exactly the same as \((-17 - x) + 2\)?

OpenStudy (anonymous):

-17 - 8 so there for x is 8

OpenStudy (anonymous):

I have no idea what you are doing.. -17 - 8 doesn't enter into it.

OpenStudy (anonymous):

i am just lost

OpenStudy (anonymous):

Ok stop doing things.. Just look at what I'm saying.

OpenStudy (anonymous):

\[-17 - x \text{ EQUALS } -6x -2\]

OpenStudy (anonymous):

okay lol i am stopping right now

OpenStudy (anonymous):

Ok? They are the same.

OpenStudy (anonymous):

That's what equals means right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so now lets use some basic arithmetic rules to clean this up a bit. First we can swap the terms on the left because addition can be rewritten in any order and it works the same.. \[-17 -x = -x -17\] right?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Do you agree that those two are equal? Is 3+5 the same as 5+3 and -3 - 5 the same as -5 - 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because you distribute right

OpenStudy (anonymous):

Commute, but yeah.

OpenStudy (anonymous):

Ok, so -17 - x is the same as -x - 17. But we also said that -17 - x was the same as -6x - 2. So that means that -x - 17 must be the same as -6x - 2.

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So now our equation looks like this: \[-x - 17 = -6x - 2\]

OpenStudy (anonymous):

that part i got is solving for x thats has me upside ways

OpenStudy (anonymous):

We'll get there.

OpenStudy (anonymous):

So if those two things are the same.. Then isn't \[6x + (-x - 17) = 6x + (-6x - 2)\]

OpenStudy (anonymous):

Because we added 6x to both sides, and if they were equal to start with, they must still be equal.

OpenStudy (anonymous):

okay thats the part that throws me off

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

its like my flipping mind goes blank and i have no idea what comes next

OpenStudy (anonymous):

Ok relax for a second and lets talk about something else first.

OpenStudy (anonymous):

or maybe i am trying too hard to understand and i get tied up

OpenStudy (anonymous):

Think of a number, any number you want. Don't tell me what it is.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now, tell me what 2 times that number is.

OpenStudy (anonymous):

184

OpenStudy (anonymous):

Ok. So now I will magically read your mind and tell you that your number is 92!

OpenStudy (anonymous):

How did I accomplish this sorcery?

OpenStudy (anonymous):

lol yes

OpenStudy (anonymous):

you divided

OpenStudy (anonymous):

But how does that help me?

OpenStudy (anonymous):

Because you told me that 2x = 182

OpenStudy (anonymous):

Well I know that if 2x = 182 then \[2x \times {1 \over 2} = 182 \times {1 \over 2}\]

OpenStudy (anonymous):

so you have to find out what 2x amounts to

OpenStudy (anonymous):

Because you said that they were the same, so anything I do to both sides will maintain equality.

OpenStudy (anonymous):

And \[2x \times \frac{1}{2} = x\]

OpenStudy (anonymous):

And \[182 \times \frac{1}{2} = 92\]

OpenStudy (anonymous):

So therefore x = 92

OpenStudy (anonymous):

So now I'm thinking of a number.

OpenStudy (anonymous):

91

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Whoops. Sorry I typo'd your response.

OpenStudy (anonymous):

Shoulda been 184 not 182.

OpenStudy (anonymous):

no one has ever broke it down like this thank you

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

so you want me to solve for your number

OpenStudy (anonymous):

Now if you take the number I'm thinking of, and add 17 to it. You will get 2 more than 6 times that same number.

OpenStudy (anonymous):

So first. can you write an equation (using x for my number) which describes the statement I've given you?

OpenStudy (anonymous):

2(6^6)+17=X

OpenStudy (anonymous):

Not quite. Lets break down what I said: x + 17 The number I'm thinking of, and add 17 to it = IS 2 more than 6 times that number 2 + 6x

OpenStudy (anonymous):

Or better still 6x + 2

OpenStudy (anonymous):

12(6)=72+17=89

OpenStudy (anonymous):

No no no.

OpenStudy (anonymous):

OH WOW

OpenStudy (anonymous):

You keep jumping ahead ;)

OpenStudy (anonymous):

okay am slowing down

OpenStudy (anonymous):

and laughing at myself

OpenStudy (anonymous):

np =)

OpenStudy (anonymous):

Try again. 14 more than the number is 2 more than 4 times the number.

OpenStudy (anonymous):

We're just practicing turning english into equations.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Don't try to solve it yet. Just rewrite what I say in equation form.

OpenStudy (anonymous):

Actually 14 more than the number is 5 more than 4 times the number.

OpenStudy (anonymous):

14 + n = 2(4)

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

Almost!

OpenStudy (anonymous):

was i close

OpenStudy (anonymous):

The left side is right, the right side is not.

OpenStudy (anonymous):

Try my last version, but break it down a bit. If x is my number what is 4 times my number??

OpenStudy (anonymous):

14 + n = 5(4 + n) oh please be right

OpenStudy (anonymous):

You're doing great btw. It's important to have patience when working through this.

OpenStudy (anonymous):

No, that's not right still. Lets work through it a little at a time.

OpenStudy (anonymous):

What is 4 times my number?

OpenStudy (anonymous):

your number is 14 right